Show that a1,a2,a3....an form an AP where an is defined as
an= 3+4n

Answer by amarjeeth123(569)

We have an=3+4n
Substituting n+1 we get an+1=3+4(n+1)=4n+4+3=4n+7
Common difference d=an+1-an=4n+7-4n-3=4
Since the common difference d is constant they are in an arithmetic progression.